OpenStudy (thesecret20111):

How would I solve: 3108416303=a^2010-a^1970?

ganeshie8 (ganeshie8):

can u tell us the source of this problem ? school/contest ?

OpenStudy (solomonzelman):

looks like a prime, although I wouldn't say for sure. You can't divide by 3, because the sum of all digits is not divisible by 3, and it is not an even number, nor it ends with a 5.

OpenStudy (thesecret20111):

I've been working on a question for school but have got stuck here. The calculator can solve this but I'm not sure where to start myself.

OpenStudy (thesecret20111):

If it is indeed a prime what does that mean as far as solving?

OpenStudy (solomonzelman):

Yeah, just verified on wolfram, it is a prime.

OpenStudy (anonymous):

you can get logarithm from that equation.

OpenStudy (solomonzelman):

No, no use, because the variables are NOT in exponents.

OpenStudy (anonymous):

log(c)=2010*log(a) - 1970*log(a)=40*log(a)=log(c) AM I RIGHT?

OpenStudy (solomonzelman):

You mean, \(\normalsize\color{black}{ 3108416303=a^{2010}-a^{1970} }\) \(\normalsize\color{black}{ \log 3108416303=\log (a^{2010}-a^{1970}) }\) ?

OpenStudy (solomonzelman):

This is even worse than before

OpenStudy (solomonzelman):

it is not log(a)-log(b), it is log(a-b)

OpenStudy (solomonzelman):

I would leave it when I say that http://www.wolframalpha.com/input/?i=3108416303%3Da%5E2010-a%5E1970.+solve+for+a.

OpenStudy (thesecret20111):

Wait so logs aren't going to help here are they?

OpenStudy (solomonzelman):

I don't think so

OpenStudy (thesecret20111):

So should I just wrote down that I used calculator to solve for a?

OpenStudy (solomonzelman):

okay, what will you do after this? you can say it is a product of 3 numbers, as a²-b²= .... because a^40= (a^20)²

OpenStudy (solomonzelman):

and product of 4 numbers and 5 number

OpenStudy (solomonzelman):

It is a quite silly problem, to my (childish) opinion

OpenStudy (thesecret20111):

A silly problem? I was simply asking if this was possible to work out and how I'd go about it.